OpenStudy (anonymous):

PLEASE HELP!!! MEDAL WILL BE REWARDED!!! Find the derivative of f(x) =-9/x at x = -8. I think it is 9/8 but I am unsure...

OpenStudy (anonymous):

set x^-1

OpenStudy (anonymous):

if you know the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) then this is easy if you have to do it by hand, which is my guess, then you have some work to do

OpenStudy (anonymous):

So its 9/64? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Oh okay thanks you so much!!!

OpenStudy (anonymous):

Would it be the opposite way around?

OpenStudy (anonymous):

oops ignore me you are right

OpenStudy (anonymous):

\(\frac{9}{64}\) is correct

OpenStudy (anonymous):

Oh okay thanks again! I have more questions im asking now!!

OpenStudy (anonymous):

i think she's right derivative would be 9/x^2 so at -8 the derivative is 9/64

OpenStudy (anonymous):

ignore me lol